A vapour pressure aqueous solution of glucose is 750nm of hg and vapour prrssure of pure water is 760nm of hg calculate the mole fraction of solute and molality of solution

Grade:12th pass

1 Answers

BIKRAM
6 years ago
This question is related to lowering of vapour pressure. Roult’s law P = X1×P0;here P= 750 nm of Hg , P0=760 nm of Hg
X1= (750/760)=0.987; Mole Fraction of solute (X2)=1-X1=0.013:
Molality = (0.013/0.987×18)×1000 = 0.73 (m)